OpenStudy (lgbasallote):

how to verify if \(e^x\) is homogeneous or not? i think it's not homogeneous because there's no y...

OpenStudy (lgbasallote):

@phi can you help me here instead? :D

OpenStudy (phi):

f(x)= e^x f(tx)= e^(tx) = (e^x)^t which is not the same as t e^x so no, it is not homogeneous

OpenStudy (lgbasallote):

but it can be turned into \[e^t + e^x\] right? does it make any difference?

OpenStudy (phi):

This can be confusing (I always think twice about these things). But no. If we replace e with 2 (for easy reckoning) we can see 2^(2*3)= 2^6= 64 2^2 + 2^3= 4+8 =12 but (2^2)^3= 4^3 = 64

OpenStudy (lgbasallote):

oops oh yeah lol it was \(e^{x+y} = e^xe^y\) lol